# Information diffusion assumptions can distort our understanding of social network dynamics

### v2

@@ Introduction
Despite significant advances in the field, data from social media platforms present important limitations (5, 26). For example, dynamic socio-technical systems continuously change, influencing user behavior in non-transparent ways even as we attempt to study them (28). The influence of evolving and opaque platform algorithms adds further complexity to analyses (29). Here we focus on an often-overlooked issue arising from the difficulty to reconstruct diffusion cascades. Some platforms, like Facebook and Instagram, only provide aggregated data about cascade sizes. Others, like Twitter (now X), Mastodon, Bluesky, and Threads, provide more data but typically obscure the pathways of information diffusion by attributing all resharing actions to the original poster. This misrepresentation conceals the true dynamics of how information spreads, hindering our ability to fully understand these processes (See Figure 1 (b, c)). To address this problem, some researchers have proposed methods to infer the structure of information cascades to predict the actual patterns of online information flow (30, 31, 32, 33, 34, 35, 36, 37). However, this problem is notoriously challenging because the reconstruction process relies on numerous assumptions, introducing significant potential for error. Additionally, this potential for error increases as the cascade grows in size, because as the number of nodes increases, it becomes increasingly difficult to accurately identify the parent of a new post (38). In rare cases, researchers may collaborate with platforms, which have access to detailed information about sharing actions (39). However, the general lack of ground-truth data renders validation of reconstruction techniques impossible (11).
@@
Because of this challenge, most studies analyze platform-provided data without reconstructing information diffusion cascades. The few studies that attempt to reconstruct the cascades typically rely on follower networks, which capture social connections between platform users and are leveraged to estimate potential content exposure (31, 33, 40). However, these approaches have not been rigorously validated. Furthermore, gathering this data is typically infeasible for researchers working with large datasets due to platform limitations.
@@
These limitations raise important concerns about the consequences of overlooking the complex dynamics of information diffusion on social media platforms. To what extent could these choices distort our understanding of online information dynamics, such as the role of influential users, the emergence of polarization, and the vulnerability of platforms and users to various forms of manipulation? In this study, we illustrate these concerns by first quantifying how bypassing the cascade reconstruction process altogether impacts measures of social influence in two case studies on Twitter and Bluesky. After determining that this omission dramatically affects assessments of node influence, we investigate the structural effects of different reconstruction approaches. Leveraging a widely studied dataset of over 100,000 Twitter news cascades (31, 41, 42, 43, 44, 45, 46), we uncover substantial discrepancies in cascades at both the micro and macro level.
@@ Results
A set of cascade trees reconstructed from the data with the PDI method can be combined into a weighted resharing network. Nodes in this network represent users and edges capture the flow of information. Specifically, a link $(i\rightarrow j,w)$ represents a directed edge from user $i$ to user $j$, weighted by $w$, the number of times user $j$ reshared user $i$’s content. However, unlike reconstruction methods that generate a single cascade in deterministic fashion (31, 33), PDI can stochastically generate many different realizations of each cascade. This allows us to construct many versions of the weighted resharing network.
@@
Pinpointing the most influential individuals within social networks is a critical and widely studied challenge across fields ranging from epidemiology (47, 48) and public health (49) to political communication (50, 51, 19) and marketing (52, 53). These key nodes can determine whether an epidemic will spread or whether a messaging campaign will achieve its intended impact.
@@
Comparing the two networks lets us determine the effects of the reconstruction method on the analyses of node influence—if the results are very similar, it would indicate that the reconstruction process has minimal impact on these analyses. To measure node influence, we calculate node out-strength, which we refer to as node strength for brevity. This is a widely recognized and intuitive metric, defined as the total number of reshares a node accumulates (54, 55, 56). As shown in Figure 2, there are important differences in node influence before and after reconstruction. In the naive resharing network, influence is concentrated among a few accounts that tend to reshare infrequently. In the reconstructed network, on the other hand, influence is more broadly distributed across many accounts, including amplifiers that tend to reshare content posted by others.
@@ Discussion
These findings were enabled by Probabilistic Diffusion Inference, a novel and flexible approach for reconstructing information cascades that allows us to fine-tune assumptions and assess their impact. The focus on the underlying assumptions makes cascade reconstruction transparent. By incorporating stochasticity, PDI also allows us to explore the variance in outcomes of interest. Moreover, the flexibility of the PDI approach allows researchers to adjust assumptions, enabling a deeper exploration of the human and algorithmic factors driving information diffusion. For example, how might diffusion dynamics shift if a platform, like X, actively promotes certain political actors, as some have suggested (59, 60, 61)? How does this differ from Meta’s new microblogging platform, Threads, which has indicated it will not insert unwanted political content into user feeds (62)? Researchers could incorporate node features, like political content, into the probability distributions to explore these and other interesting research questions.
@@
Despite these benefits, PDI should not be considered “more accurate” than other techniques, such as the Time-Inferred Diffusion. The validation of inference methods requires the availability of ground-truth information diffusion data (63, 64, 65). If platforms were to publicly share cascade data with researchers, the PDI framework could be leveraged to refine assumptions and optimize parameter settings for more accurate modeling. However, we note that even platforms have to make assumptions about parent attribution, as users may be exposed to a piece of content in different ways prior to sharing, which are not revealed by their specific sharing action.
@@
The substantial divergence we have found between networks reconstructed with different methods underscores the potential risks for researchers who study online phenomena—especially when relying on naive networks provided by platforms. Given the widespread reliance on platform-provided data and the lack of ground truth for diffusion cascades, researchers must approach these analyses with caution. Such data is inherently complex and may be incomplete, exacerbating the challenge of accurately capturing underlying dynamics. Therefore, it is essential to develop methods that can effectively account for these limitations. These issues have far-reaching implications for fields that analyze social media networks, such as conservation science (66), political communication (50, 51, 19), public health (49), and epidemiology (67, 68). Future research should focus on identifying which analyses are most sensitive to reconstruction methods and ensuring their robustness across varying assumptions.
@@
Computational social science must continue to develop innovative analytical approaches that make transparent assumptions and are robust to rigorous methodological scrutiny, transparency in assumptions (69, 70, 71). Such progress is crucial for deepening our understanding of complex digital ecosystems and the social dynamics that unfold within them.
@@ Methods
Bluesky is a decentralized, Twitter-like micro-blogging platform designed to offer a federated social experience (73, 74). We collected all data from Bluesky between March 1, 2024, and March 14, 2024, using the public Firehose endpoint, which streams all posts shared on the platform (75). We then randomly sampled 5,000 repost cascades originating in the first seven days of this period, following the platform’s public launch (76). The same sampling procedure used for the U.S. Midterm dataset was applied, capturing reposts up to one week later (March 21, 2024). We excluded 290 cascades from our analysis: 271 due to missing metadata for at least one user’s follower count, and 19 because of timestamp discrepancies caused by Bluesky’s distributed architecture (77). This resulted in a final dataset of 4,710 cascades consisting of 21,338 posts from 15,550 users.
@@
PDI enables flexible computation of the probabilities $p_{ij}$ using researcher-defined assumptions. In this work, we adopt two common assumptions. First, users with more followers are more likely to be the parents of a resharing user (78), which we refer to as the followers assumption. Second, users who recently reshared the content are more likely to be the true parents of subsequent users (79, 80), referred to as the recency assumption.
@@
The recency assumption is modeled using a power-law distribution, which has been shown to describe the timing of resharing behavior on social media platforms (79, 80):
@@ Acknowledgments
We are grateful to Alessandro Flammini and Jonas Juul for their valuable discussions and feedback. We also thank Soroush Vosoughi, Deb Roy, and Sinan Aral for generously providing the data used in our information cascade structure analyses. This work used JetStream2 (81, 82) at Indiana University through allocation CIS200033 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296. We also acknowledge the team of developers at the Observatory on Social Media for their assistance in collecting Bluesky data.

### v3

@@ Introduction
Despite significant advances in the field, data from social media platforms present important limitations (5, 26). For example, dynamic socio-technical systems continuously change, influencing user behavior in non-transparent ways even as we attempt to study them (28). The influence of evolving and opaque platform algorithms adds further complexity to analyses (29). Here we focus on an often-overlooked issue arising from the difficulty to reconstruct diffusion cascades. Some platforms, like Facebook and Instagram, only provide aggregated data about cascade sizes. Others, like Twitter (now X), Mastodon, Bluesky, and Threads, provide more data but typically obscure the pathways of information diffusion by attributing all resharing actions to the original poster. This misrepresentation conceals the true dynamics of how information spreads, hindering our ability to fully understand these processes (See Figure 1 (b, c)). To address this problem, some researchers have proposed methods to infer the structure of information cascades to predict the actual patterns of online information flow (30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41). However, this problem is notoriously challenging because the reconstruction process relies on numerous assumptions, introducing significant potential for error. Additionally, this potential for error increases as the cascade grows in size, because as the number of nodes increases, it becomes increasingly difficult to accurately identify the parent of a new post (42). In rare cases, researchers may collaborate with platforms, which have access to detailed information about sharing actions (43, 33, 34). However, the general lack of ground-truth data renders validation of reconstruction techniques impossible (11).
@@
Because of this challenge, most studies analyze platform-provided data without reconstructing information diffusion cascades. The few studies that attempt to reconstruct the cascades typically rely on follower networks, which capture social connections between platform users and are leveraged to estimate potential content exposure (30, 44, 38, 40). However, these approaches have not been rigorously validated. Furthermore, gathering this data is typically infeasible for researchers working with large datasets due to platform limitations.
@@
These limitations raise important concerns about the consequences of overlooking the complex dynamics of information diffusion on social media platforms. To what extent could these choices distort our understanding of online information dynamics, such as the role of influential users, the emergence of polarization, and the vulnerability of platforms and users to various forms of manipulation? In this study, we illustrate these concerns by first quantifying how bypassing the cascade reconstruction process altogether impacts measures of social influence in two case studies on Twitter and Bluesky. After determining that this omission dramatically affects assessments of node influence, we investigate the structural effects of different reconstruction approaches. Leveraging a widely studied dataset of over 100,000 Twitter news cascades (40, 45, 46, 47, 48, 49, 50), we uncover substantial discrepancies in cascades at both the micro and macro level.
@@ Results
A set of cascade trees reconstructed from the data with the PDI method can be combined into a weighted resharing network. Nodes in this network represent users and edges capture the flow of information. Specifically, a link $(i\rightarrow j,w)$ represents a directed edge from user $i$ to user $j$, weighted by $w$, the number of times user $j$ reshared user $i$’s content. However, unlike reconstruction methods that generate a single cascade in deterministic fashion (40, 38), PDI can stochastically generate many different realizations of each cascade. This allows us to construct many versions of the weighted resharing network.
@@
Pinpointing the most influential individuals within social networks is a critical and widely studied challenge across fields ranging from epidemiology (51, 52) and public health (53) to political communication (54, 55, 19) and marketing (56, 57, 30). These key nodes can determine whether an epidemic will spread or whether a messaging campaign will achieve its intended impact.
@@
Comparing the two networks lets us determine the effects of the reconstruction method on the analyses of node influence—if the results are very similar, it would indicate that the reconstruction process has minimal impact on these analyses. To measure node influence, we calculate node out-strength, which we refer to as node strength for brevity. This is a widely recognized and intuitive metric, defined as the total number of reshares a node accumulates (58, 59, 60). As shown in Figure 2, there are important differences in node influence before and after reconstruction. In the naive resharing network, influence is concentrated among a few accounts that tend to reshare infrequently. In the reconstructed network, on the other hand, influence is more broadly distributed across many accounts, including amplifiers that tend to reshare content posted by others.
@@ Discussion
These findings were enabled by Probabilistic Diffusion Inference, a novel and flexible approach for reconstructing information cascades. By combining stochastically generated realizations of each cascade, we are able to construct many versions of weighted resharing (influence) networks. This allows us to explore the variance in outcomes of interest. An extension that we have not explored in this manuscript is probabilistically reflecting the influence of multiple parents on a node within a single cascade. By combining multiple reconstruction realizations, we can represent a cascade as a probability-weighted acyclic graph rather than a simple tree. This would make it possible to causally attribute a reshare action not only to one previous action, but to multiple prior exposures (33, 34).
@@
The focus of PDI on the underlying assumptions also makes cascade reconstruction transparent. This helps researchers fine-tune assumptions and assess their impact, enabling a deeper exploration of the human and algorithmic factors driving information diffusion. For example, how might diffusion dynamics shift if a platform, like X, actively promotes certain political actors, as some have suggested (63, 64, 65)? How does this differ from Meta’s new microblogging platform, Threads, which has indicated it will not insert unwanted political content into user feeds (66)? Researchers could incorporate node features, like political content, into the probability distributions to explore these and other interesting research questions.
@@
Despite these benefits, PDI should not be considered “more accurate” than other techniques, such as the Time-Inferred Diffusion. The validation of inference methods requires the availability of ground-truth information diffusion data (67, 68, 69). If platforms were to publicly share cascade data with researchers, the PDI framework could be leveraged to refine assumptions and optimize parameter settings for more accurate modeling. However, we note that even platforms have to make assumptions about parent attribution, as users may be exposed to a piece of content in different ways prior to sharing, which are not revealed by their specific sharing action.
@@
The substantial divergence we have found between networks reconstructed with different methods underscores the potential risks for researchers who study online phenomena—especially when relying on naive networks provided by platforms. Given the widespread reliance on platform-provided data and the lack of ground truth for diffusion cascades, researchers must approach these analyses with caution. Such data is inherently complex and may be incomplete, exacerbating the challenge of accurately capturing underlying dynamics. Therefore, it is essential to develop methods that can effectively account for these limitations. These issues have far-reaching implications for fields that analyze social media networks, such as conservation science (70), political communication (54, 55, 19), public health (53), and epidemiology (71, 72). Future research should focus on identifying which analyses are most sensitive to reconstruction methods and ensuring their robustness across varying assumptions.
@@
Computational social science must continue to develop innovative analytical approaches that make transparent assumptions and are robust to rigorous methodological scrutiny, transparency in assumptions (73, 74, 75). Such progress is crucial for deepening our understanding of complex digital ecosystems and the social dynamics that unfold within them.
@@ Methods
Bluesky is a decentralized, Twitter-like micro-blogging platform designed to offer a federated social experience (77, 78). We collected all data from Bluesky between March 1, 2024, and March 14, 2024, using the public Firehose endpoint, which streams all posts shared on the platform (79). We then randomly sampled 5,000 repost cascades originating in the first seven days of this period, following the platform’s public launch (80). The same sampling procedure used for the U.S. Midterm dataset was applied, capturing reposts up to one week later (March 21, 2024). We excluded 290 cascades from our analysis: 271 due to missing metadata for at least one user’s follower count, and 19 because of timestamp discrepancies caused by Bluesky’s distributed architecture (81). This resulted in a final dataset of 4,710 cascades consisting of 21,338 posts from 15,550 users.
@@
PDI enables flexible computation of the probabilities $p_{ij}$ using researcher-defined assumptions. In this work, we adopt two common assumptions. First, users with more followers are more likely to be the parents of a resharing user (82), which we refer to as the followers assumption. Second, users who recently reshared the content are more likely to be the true parents of subsequent users (83, 84), referred to as the recency assumption.
@@
The recency assumption is modeled using a power-law distribution, which has been shown to describe the timing of resharing behavior on social media platforms (83, 84):
@@ Acknowledgments
We are grateful to Alessandro Flammini, Jonas Juul, and Dean Eckles for their valuable feedback. We also thank Soroush Vosoughi, Deb Roy, and Sinan Aral for generously providing the data used in our information cascade structure analyses. This work used JetStream2 (85, 86) at Indiana University through allocation CIS200033 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296. We also acknowledge the team of developers at the Observatory on Social Media for their assistance in collecting Bluesky data.